OpenStudy (anonymous):

evaluate f(2) for the function f(x). f(x)=5x^2+3x-6 1.32 2.112 3.8 4.20

OpenStudy (anonymous):

Just sub x for 2

OpenStudy (anonymous):

\[ f(2)=5(2)^2+3(2)-6 \]

OpenStudy (anonymous):

5 times 2 is 10

OpenStudy (anonymous):

soo \[ f(2)=10^2+3(2)-6 \]

OpenStudy (anonymous):

10^2 = 100

OpenStudy (anonymous):

so \[ f(2)=100+3(2)-6 \]

OpenStudy (anonymous):

then 3 times 2 is 6 and you subract 6 so... \[f(2)=100\]

OpenStudy (anonymous):

:)

OpenStudy (mrnood):

It's 5*(2)^2 NOT (5*2)^2

OpenStudy (mrnood):

\[5x ^{2}+3x-6\] Substitute x=2 gives \[5(2)^{2}+3(2)-6\] Solve that for your answer

OpenStudy (anonymous):

Whoops